A man travels for Town X to Town Y at an average rate of 50 mph and returns at an average rate of 40 mph. He takes 1/2 hr. longer than he would take if he made the round trip at an average of 45 mph. What is the distance from Town X to Town Y?
Answer by checkley75(3666)

d/50+d/40=(2d/45)+.5
(40d+50d)/2000=(2d+22.5)/45
90d/2000=(2d+22.5)/45 now cross multiply.
45*90d=4000d+45000
4050d-4000d=45000
50d=45000
d=45000/50
d=900 miles for the trip fron x to y.
proof
900/50+900/40=1800/45+.5
18+22.5=40+.5
40.5=40.5